Three kilograms of an ideal gas with R = 204 J/kg-K and k = 1.667 have 316 kJ of heat added during an isobaric process. The initial temperature is 38oC. Determine (a) Wnf, (b) U, (c) S.


1
Expert's answer
2021-12-06T09:40:32-0500

"C_p=\frac{kR}{k-1}=510~\frac{J}{kg\cdot K},"

"T_2=T_1+\frac Q{mC_p}=518~K,"

"C_V=\frac{R}{k-1}=306~\frac{J}{kg\cdot K},"

"\nu=\frac{C_p-C_V}{R}=1~mol,"

a)

"A=\nu R\Delta T=42~kJ,"

b)

"U=Q-A=274~kJ,"

c)

"S=mC_p\ln\frac{T_2}{T_1}=780~\frac JK."